5/22-23: MZ release from PSL
5/20-24: Exchanging cryo valves.
- 5/27-6/2: Y arm mass lock.
5/29: RT PC maintenance -> go to safe mode for all VIS connected Dolphin (specially BS)
- 6/3-6/14: Installation of TMS for ETMY
- 6/8: Engineering run
5/22AM~: tuned off cryostat cooler, finesse measurement for Xarm
5/22AM~: tuned off shield cooler, finesse measurement for Xarm
5/2: Cooling start
5/23AM~: tuned off cryostat cooler, finesse measurement for Yarm
5/17~: tuned off ETMY cryostat cooler
- Long term:
- 4/22-5/31: Yarm commissioning
- 5/27-6/30: DRMI commissioning
- 7/1-8/9: DRFPMI commissioning
- 8/12-10/4: DRFPMI ASC
- 10/7-11/15: Post Commissioning
- 11/18-12/13: Noise hunting
Report of work
- Type A: VIS team is changing all the TYPE A models for damping schemes.
- Type A, CRY, COM: ITMX is fluctuating a lot.
- COM: Yarm finesse was 1290, a bit lower than before. It might be contamination.
- VIS damping worked well.
- COM: Y arm lock: New filters will be applied tonight.
- Killing ITMX resonance.
- Finesse measurement for both arm.